# An evaluation of LLMs and Google Translate for translation of selected Indian languages via sentiment and semantic analyses

### v2

@@ 4.3 Semantic Analysis
Table 10: Chapter-wise cosine similarity between language models and Human Expert for Sanskrit-English translation.
@@
Table 11: Semantic analysis showing cosine similarity scores (Sanskrit-English translation).
@@
Table 12: Semantic analysis showing cosine similarity scores (Telugu-English translation).
@@
Table 13: Semantic analysis showing cosine similarity scores (Hindi-English translation).)
We further evaluate the accuracy of Hindi and Telugu translations, with the mean and standard deviation of the similarity scores. We grouped the sentences into bins of 10 sentences to calculate the mean similarity score and standard deviation.
@@
In the case of Telugu-English translation ( Figure 8), the similarity scores declined significantly across all models for sentences in the 10–40 range. These sentences contained extensive use of personification, imagery, symbolism, and abstract elements, which posed challenges for translation. Examples of such sentences include: “Roar forth, oh ocean of flames!” and “The mountain moves like a massive chariot!” The complexity of these expressions contributed to the lower similarity scores. Despite this, GPT-4o and Gemini remained the most reliable models, whereas GPT-3.5 exhibited significant inconsistencies.
@@
In the case of Hindi-English translation ( Figure 9), we observe that sentences in the 1–10 and 90–97 ranges exhibited higher similarity to human translations. A closer analysis of the translated texts revealed that these sentences primarily consisted of clear and direct narration with minimal use of figurative language. Furthermore, sentences in the 90–97 range demonstrated the least variability, as reflected in Figure 9) for similar reasons. Among the models, GPT-4o and Gemini consistently outperformed others when compared to Human Expert translation.
@@
4.4 Qualitative analysis of semantically similar verses
@@
We qualitatively evaluate how well the translation has been done for the semantically most similar verses and least similar verses to give us an insight into how well the models are performing and whether they correctly transmit the original meaning of the text
@@
As shown in Table 14, in general, we find that the respective language models show high accuracy in translating structured, fact-based Sanskrit (Bhagavad Gita) statements while struggling with abstract, poetic, and philosophical content. Sentences with direct relationships and clear syntactic patterns, such as enumerations (Chapter 10, verse 31): ”Of the purifiers, I am the wind; of the wielders of weapons, I am Rama”), achieve the highest similarity scores (close to 0.90). In contrast, abstract and philosophical sentences discussing concepts like the imperishability of the soul or renunciation score significantly lower (around 0.10–0.40) due to their metaphorical language and cultural references. The emotional and metaphoric language expressions bring further challenges to language models with verses such as Arjuna’s distress in battle (Chapter 1, verse 28): ”my limbs fail and my mouth is drying up,” receive low similarity scores (0.36 for GPT-4o, 0.37 for GPT-3.5), as models struggle to convey the intended sentiment. Similarly, honorifics and formal expressions are often inconsistently translated, with phrases such as ”O best of the Indians” being altered, affecting both tone and cultural accuracy. These inconsistencies highlight the limitations of current language models in handling context-dependent translations. Furthermore, GPT-4o and GPT-3.5 perform well on structured statements but struggle with literary and philosophical Sanskrit texts (Chapter 2, verse 30).
@@
In the case of Telugu-English translation (Table 15, sentences with a clear structure, like those describing a lamp turning into a star, scored the highest across all models. GPT-4o, GPT-3.5, and Gemini all had similarity scores above 0.9, indicating they effectively captured the sequence of events. Minor differences in phrasing indicate variations in sentence construction rather than a loss of meaning, as seen in verse 48. The Human Expert translation states: ”The star sang its inviting song. The lamp went out. It transformed into a star.” Meanwhile, the Google translation reads: ”The celestial star sang an invocation. The lamp went out. Turned into a star.” Despite differences in wording, the overall meaning remains unchanged. The descriptive and philosophical sentences, such as those discussing the vastness of the sky or the feeling of disappearing into the cosmos as described in sentences 33 and 14, respectively. They showed moderate similarity (0.354–0.520), which relies on figurative language, which the models often interpret differently, leading to minor shifts in meaning. GPT-4o was consistent in maintaining the original tone, though some rewording was observed.
@@
In the case of Hindi-English translation (Tamas) shown in Table 16, structured and action-oriented sentences show the highest similarity, with GPT-4o scoring 0.927. However, figurative and descriptive elements resulted in lower scores. For example (Sentence 12), in ”thorny bristles on its back,” the models struggled to retain details, leading to scores of 0.520 (GPT-4o), 0.317 (GPT-3.5), and 0.364 (Gemini). Similarly, metaphorical expressions such as ”he cursed the evil moment” had lower scores: 0.536 (GPT-4o), 0.535 (GPT-3.5), and 0.498 (Gemini). The sensory descriptions, such as the pig’s movement, showed moderate similarity (0.919 for GPT-4o), suggesting that while models captured the overall meaning, they missed finer details. Emotionally charged phrases, such as ”the dagger in his hand looked odd and irrelevant,” showed greater divergence, with scores of 0.785 (GPT-4o), 0.655 (GPT-3.5), and 0.700 (Gemini).
@@
Table 14: Selected most and least semantic similar verses from the Bhagavad Gita (Sanskrit-English) translation.
@@
Table 15: Selected most and least semantic similar verses from the Maha Prashatham (Telugu-English translation).
@@
Table 16: Selected most and least semantic similar verses from Tamas (Hindi-English) translation.
5 Discussion
We first summarise the results from the different investigations that include sentiment and semantic analysis for comparing the respective translations of the given texts (Bhagavad Gita, Tamas, and Maha P). The sentiment analysis results show that pessimistic, anxious, and sad tones appear across all models in Sanskrit translations, reflecting the solemn and philosophical nature of the Bhagavad Gita. However, sentiment distribution varied significantly between models (Table x). In general, we found that Google Translate presented the highest extremes in sentiment polarity, with both the highest optimistic and pessimistic sentiment counts, indicating an inconsistency in emotional tone (Table x). In contrast, GPT-4o closely aligns with human translations, maintaining a more balanced sentiment distribution.
Considering the Hindi-English translation of Tamas, the Human Expert version had the most pessimistic sentiments, indicating a stronger focus on negative tones (Figure - Table x). GPT-4o showed the highest Anxious sentiment, suggesting it may amplify emotional intensity. Google Translate had the most Sad and Annoyed sentiments, reflecting a generally negative tone. Gemini had the highest Optimistic sentiment, making its translations more positive in comparison.
@@ 5 Discussion
In Telugu translations, sentiment distribution is relatively more consistent across models. However, Google Translate introduces more emotional variation, generating translations that show empathy and denial. GPT-4o and Gemini again lean toward pessimism, similar to their Hindi translations. This consistency suggests that while LLMs can recognise broad emotional trends, they still exhibit model-specific biases in sentiment interpretation. The sentiment analysis highlights the importance of context in translation. Many variations suggest that LLMs translate text into English without integrating background knowledge from the source material. Providing contextual prompts could enhance translation quality, ensuring the emotional tone aligns more closely with the original text. Further research could explore whether incorporating historical, cultural, or philosophical context improves sentiment preservation.
@@
The semantic similarity analysis across Sanskrit, Telugu, and Hindi shows varying model performances. In Sanskrit, GPT-4o achieves the highest average similarity and maintains consistency, while GPT-3.5 reaches the highest peak similarity but fluctuates more (Table, Figure x). Google Translate performs the weakest, struggling with Sanskrit’s complex syntax and poetic structure. All models translate structured, fact-based Sanskrit well but struggle with abstract and philosophical content. Sentences discussing deep concepts, such as the imperishability of the soul, show lower similarity scores, indicating that metaphorical and figurative language remains a challenge.
@@
In the case of Telugu-English translation, Gemini outperforms other models in overall semantic similarity, followed by GPT-4o. GPT-3.5 scores significantly lower, indicating weaker alignment with human translations (Table x). Structured sentences describing sequences of events, such as a lamp turning into a star, achieve high similarity scores, suggesting that LLMs excel at preserving factual meaning. However, figurative expressions about cosmic vastness or abstract emotions show moderate similarity, implying a tendency for models to rephrase rather than directly translate poetic language.
@@
The evaluation of Hindi translations across different models using cosine similarity scores shows that GPT-4o provides the most accurate results (Table, Figure x), closely matching human translations. It maintains the meaning, sentence structure, and context of the original text better than other models. Gemini also performs well but sometimes lacks the level of accuracy seen in GPT-4o. GPT-3.5 provides moderate accuracy but tends to simplify certain phrases, which can lead to minor changes in meaning. Google Translate has the lowest accuracy, often struggling with sentence structure and context. The results suggest that as language models improve, they better capture the meaning of the text, with GPT-4o currently being the most reliable for Hindi translation. Overall, GPT-4o emerged as the most reliable model across all three languages, excelling in Sanskrit and Hindi (Table x). Gemini performs well in Telugu but struggles with Sanskrit and Hindi, indicating it may be optimised for specific linguistic structures. GPT-3.5 shows strong peak similarity in Sanskrit but lacks overall consistency. Google Translate remains the weakest performer, particularly in Sanskrit and Hindi, highlighting its limitations in handling complex linguistic structures.
@@
In terms of the limitations, we note that we used translation by one human expert to compare LLMs and Google Translate. The texts such as the Bhagavad Gita have been translated hundreds of times  in the last two millenniums, and we note that even the human expert translations have variations when it comes to vocabulary, sentiments and semantics as shown in earlier studies [65]. Furthermore, texts such as the Bhagavad Gita feature the philosophical foundation of Hinduism and are full of allegories and metaphors, which have cultural and religious contexts, making them harder to translate. The composition of the text was more than 2500 years ago and orally transmitted over a millennium before being written down. Therefore, it is not an easy test for translation, both for humans and language models. Taking these into account, we considered Hindi which is the largest language (native speaker) in India and Telugu which has over 83 million native speakers (based on the 2011 Indian census) [120]. We have only considered one text from these languages and the texts we selected are classical texts (Tamas and Maha Prasthanam) which is a limitation. Further work needs to be done taking modern texts and more recent works into account, including newspapers and social media as these are more prominently auto-translated by language models. We also need to consider more human expert translations for effectively evaluating the strength of novel language models for machine translation.
@@
Future research could explore improving translations by incorporating contextual background, cultural cues, and domain-specific fine-tuning to enhance both sentiment alignment and semantic accuracy. LLMs offer more advantages over Google Translate and Human Expert since the user can promote LLMs to provide a specific vocabulary style and tone in terms of sentiments. Furthermore, the translated text by human experts can also be transformed by LLMs to get different sets of vocabulary, for example, if the translation of the Bhagavad Gita is from the 1920s, LLMs can be used to give modern rendering in terms of vocabulary which would be easier to understand by modern readers. Our framework is general and has been used in previous works where the Bhagavad Gita human expert translation has been compared [65], along with the effectiveness of Google Translate [74]. This study motivates future work for the extension of the framework for the evaluation of other languages, particularly high-resource languages for effective evaluation as more resources and data are available, such as the case of European languages (such as Spanish, French, and Italian). This would better evaluate the cross-cultural strengths and limitations of LLMs in language translation. The framework can also be extended to study how certain languages influence and are coupled with each other, and how certain metaphors and linguistic expressions overlap. Examples include Urdu vs Hindi and Hindi vs Hinglish which is an evolving language of the internet and prominent in social media platforms such as X (Twitter). In summary, our study and framework have paved the way ahead for the automated evaluation of language translation models.
6 Conclusion

### v3

@@ IV-C Semantic Analysis
TABLE X: Chapter-wise cosine similarity between language models and Human Expert for Sanskrit-English translation.
@@
TABLE XI: Semantic analysis showing cosine similarity scores (Sanskrit-English translation).
@@
TABLE XII: Semantic analysis showing cosine similarity scores (Telugu-English translation).
@@
TABLE XIII: Semantic analysis showing cosine similarity scores (Hindi-English translation).
We further evaluate the accuracy of Hindi and Telugu translations, with the mean and standard deviation of the similarity scores. We group the sentences into bins of 10 sentences to calculate the mean similarity score and standard deviation.
@@
In the case of Telugu-English translation ( Figure 8), we find that the the similarity scores declined significantly across all models for sentences in the 10–40 range. These sentences contained extensive use of personification, imagery, symbolism, and abstract elements, which posed challenges for translation. Examples of such sentences include: “Roar forth, oh ocean of flames!” and “The mountain moves like a massive chariot!” The complexity of these expressions contributed to the lower similarity scores. Despite this, GPT-4o and Gemini remained the most reliable models, whereas GPT-3.5 exhibited significant inconsistencies.
@@
In the case of Hindi-English translation ( Figure 9), we observe that sentences in the 1–10 and 90–97 ranges exhibit higher similarity to human translations. A closer analysis of the translated texts revealed that these sentences primarily consist of clear and direct narration with minimal use of figurative language. Furthermore, the sentences in the 90–97 range demonstrate the least variability, as reflected in Figure 9) for similar reasons. Among the models, GPT-4o and Gemini consistently outperform others when compared to Human Expert translation.
@@
IV-D Qualitative analysis of semantically similar verses
@@
We qualitatively evaluate how well the translation has been done for the semantically most similar verses and least similar verses to give us an insight into how well the models are performing and whether they correctly transmit the original meaning of the text.
@@
As shown in Table XIV, in general, we find that the respective language models show high accuracy in translating structured, fact-based Sanskrit (Bhagavad Gita) statements, while struggling with abstract, poetic, and philosophical content. We find that the sentences with direct relationships and clear syntactic patterns, such as enumerations (Chapter 10, Verse 31): ”Of the purifiers, I am the wind; of the wielders of weapons, I am Rama”), achieve the highest similarity scores (close to 0.90). In contrast, abstract and philosophical sentences discussing concepts such as the imperishability of the soul or renunciation score significantly lower (around 0.10–0.40) due to their metaphorical language and cultural references. The emotional and metaphoric language expressions bring further challenges to language models with verses such as Arjuna’s distress in battle (Chapter 1, verse 28): ”my limbs fail and my mouth is drying up,” receive low similarity scores (0.36 for GPT-4o, 0.37 for GPT-3.5), as models struggle to convey the intended sentiment. Similarly, we find that honorifics and formal expressions are often inconsistently translated, with phrases such as ”O best of the Indians” being altered, affecting both tone and cultural accuracy. These inconsistencies highlight the limitations of current language models in handling context-dependent translations. Furthermore, GPT-4o and GPT-3.5 perform well on structured statements but struggle with literary and philosophical Sanskrit texts (Chapter 2, verse 30).
@@
In the case of Telugu-English translation (Table XV, sentences with a clear structure, like those describing a lamp turning into a star, scored the highest across all models. GPT-4o, GPT-3.5, and Gemini all had similarity scores above 0.9, indicating they effectively captured the sequence of events. The minor differences in phrasing indicate variations in sentence construction rather than a loss of meaning, as seen in verse 48. The Human Expert translation states: ”The star sang its inviting song. The lamp went out. It transformed into a star.” Meanwhile, the Google translation reads: ”The celestial star sang an invocation. The lamp went out. Turned into a star.” Despite differences in wording, the overall meaning remains unchanged. The descriptive and philosophical sentences, such as those discussing the vastness of the sky or the feeling of disappearing into the cosmos, as described in sentences 33 and 14, respectively. They showed moderate similarity (0.354–0.520), which relies on figurative language, which the models often interpret differently, leading to minor shifts in meaning. GPT-4o was consistent in maintaining the original tone, though some rewording was observed.
@@
In the case of Hindi-English translation (Tamas) shown in Table XVI, structured and action-oriented sentences show the highest similarity, with GPT-4o scoring 0.927. However, figurative and descriptive elements resulted in lower scores. For example (Sentence 12), in ”thorny bristles on its back,” the models struggled to retain details, leading to scores of 0.520 (GPT-4o), 0.317 (GPT-3.5), and 0.364 (Gemini). Similarly, metaphorical expressions such as ”he cursed the evil moment” had lower scores: 0.536 (GPT-4o), 0.535 (GPT-3.5), and 0.498 (Gemini). The sensory descriptions, such as the pig’s movement, showed moderate similarity (0.919 for GPT-4o), suggesting that while models captured the overall meaning, they missed finer details. Emotionally charged phrases, such as ”the dagger in his hand looked odd and irrelevant,” showed greater divergence, with scores of 0.785 (GPT-4o), 0.655 (GPT-3.5), and 0.700 (Gemini).
@@
TABLE XIV: Selected most and least semantically similar verses from the Bhagavad Gita (Sanskrit-English) translation.
@@
TABLE XV: Selected most and least semantic similar verses from the Maha Prashatham (Telugu-English translation).
@@
TABLE XVI: Selected most and least semantically similar verses from Tamas (Hindi-English) translation.
V Discussion
We first summarise the results from the different investigations that include sentiment and semantic analysis for comparing the respective translations of the given texts (Bhagavad Gita, Tamas, and Maha Prasthanan). The sentiment analysis results show that pessimistic, anxious, and sad tones appear across all models in the Sanskrit-English translation (Figure 2), reflecting the solemn and philosophical nature of the Bhagavad Gita. However, sentiment distribution varied significantly between models (Figures 6 and 7). GPT-3.5 reported the closest sentiment polarity scores when compared to the human expert (Table 9). Considering the Hindi-English translation of Tamas, the Human Expert version had the most pessimistic sentiments, indicating a stronger focus on negative tones (Figure 4). GPT-4o showed the highest Anxious sentiment, suggesting it may amplify emotional intensity. Google Translate had the most Sad and Annoyed sentiments, reflecting a generally negative tone, but also expressed the most empathy when compared to Human expert and other models. Gemini had the highest Optimistic sentiment, making its translations more positive in comparison. In Telugu translation (Figure 3), sentiment distribution is relatively more consistent across models. However, Google Translate introduced more emotional variation, generating translations that show empathy and denial. GPT-4o and Google Translate leaned toward pessimism, similar to their Hindi translations. This consistency suggests that while LLMs can recognise broad emotional trends, they still exhibit model-specific biases in sentiment interpretation. In general, all models have been close to the Human expert for most of the sentiments, except for empathetic sentiments in Telugu-English (Figure 3) and Hindi-English (Figure 4) translations. These models have captured the Empathetic sentiment better in the case of the Sanskrit-English translation (Figure 2). The difference in performance could be due to the nature of the texts, i.e more empathetic sentiments were expressed in the Bhagavad Gita when compared to Tamas and Maha Prasthanam. Furthermore, we have a larger text data size in the case of the Bhagavad Gita, as all 18 chapters were considered when compared to selected chapters from Tamas and Maha Prasthanam.
@@ V Discussion
The sentiment analysis highlights the importance of context in translation. Many variations suggest that LLMs translate text into English without integrating background knowledge from the source material. Providing contextual prompts could enhance translation quality, ensuring the emotional tone aligns more closely with the original text. Further research could explore whether incorporating historical, cultural, or philosophical context improves sentiment preservation.
@@
The semantic similarity analysis across Sanskrit, Telugu, and Hindi gave varying model performances (Table 10). In Sanskrit, GPT-4o achieved the highest average similarity and maintained consistency, while Google Translate struggled with Sanskrit’s complex syntax and poetic structure. We observed that all the respective models translated structured, fact-based Sanskrit phrases well, but struggled with abstract and philosophical content featuring metaphors. The sentences discussing deep concepts, such as the imperishability of the soul (Table 14), produced lower similarity scores, indicating that metaphorical and figurative language remains a challenge.
@@
In the case of Telugu-English translation (Figure 8), Gemini outperformed other models in overall semantic similarity, followed by GPT-4o. GPT-3.5 scores significantly lower, indicating weaker alignment with human translations. Structured sentences describing sequences of events, such as a lamp turning into a star, achieve high similarity scores, suggesting that LLMs excel at preserving factual meaning. However, figurative expressions about cosmic vastness or abstract emotions show moderate similarity, implying a tendency for models to rephrase rather than directly translate poetic language.
@@
The evaluation of Hindi translations across different models using cosine similarity scores shows that GPT-4o provides the most accurate results (Figure 9), closely matching human translations. It maintains the meaning, sentence structure, and context of the original text better than other models. Gemini also performs well, but sometimes lacks the level of accuracy seen in GPT-4o. GPT-3.5 provides moderate accuracy but tends to simplify certain phrases, which can lead to minor changes in meaning. Google Translate has the lowest accuracy, often struggling with sentence structure and context. The results suggest that as language models improve, they better capture the meaning of the text, with GPT-4o currently being the most reliable for Hindi translation. Overall, GPT-4o emerged as the most reliable model across all three languages, excelling in Sanskrit (Table 10), and Telugu and Hindi translations (Figures 8 and 9). Gemini performs well in Telugu but struggles with Sanskrit and Hindi, indicating it may be optimised for specific linguistic structures. GPT-3.5 shows strong peak similarity in Sanskrit but lacks overall consistency. Google Translate remains the weakest performer, particularly in Sanskrit and Hindi, highlighting its limitations in handling complex linguistic structures. However, we note that if we take the standard deviation into consideration (Table 10 and Figures 8 and 9), across all results, no one model is significantly better than another.
@@
In terms of the limitations, we note that we used translation by one human expert to compare LLMs and Google Translate. The texts such as the Bhagavad Gita have been translated hundreds of times  in the last two millennia, and we note that even the human expert translations have variations when it comes to vocabulary, sentiments and semantics as shown in earlier studies [65]. Furthermore, texts such as the Bhagavad Gita feature the philosophical foundation of Hinduism and are full of allegories and metaphors, which have cultural and religious contexts, making them harder to translate. The composition of the text was more than 2500 years ago and orally transmitted over a millennium before being written down. Therefore, it is not an easy test for translation, both for humans and language models. Taking these into account, we considered Hindi, which is the largest language (native speaker) in India and Telugu which has over 83 million native speakers (based on the 2011 Indian census) [121]. We have only considered one text from these languages, and the texts we selected are classical texts (Tamas and Maha Prasthanam), which is a limitation. Furthermore, we selected only one translation per text since we reviewed three different texts and three LLMs, additional translations can be added in future work. Further work needs to be done, taking modern texts and more recent works into account, including newspapers and social media, as these are more prominently auto-translated by language models. We also need to consider more human expert translations for effectively evaluating the strength of novel language models for machine translation. Since LLMs are black box models, interpretability of how decision was made is a challenging task which depends on model architecture, training strategy and data [122]. The way the LLM makes a decision depends mostly on the model architecture and data used in pre-trained Transformer-based models. The training data for the respective LLMs are different, and hence they make slightly different decisions, for example, our results show that GPT-4o maintains emotions better in translations of philosophical texts such as the Bhagavad Gita.
It is difficult even for human translators to maintain cultural differences and differences in language structure to maintain emotional and semantic retention in translation. Previous study has shown that there is a large difference in the use of vocabulary by human translations of philosophical texts such as the Bhagavad Gita [65]. However, in the same study, it has also been found that the semantic and emotional similarity were maintained despite the difference in vocabulary. We note that different periods in time influence the vocabulary of spoken and written language; therefore, the results depends on what periods (writing style and vocabulary) contributed to the majority of the training data corpus for the LLM.
The results from this study about the assessment of LLMs can guide future research in the area of language translation for low-resource languages. The observations from this study, including the strengths and weaknesses of the respective models can be used as a guideline to build better models. Better translation models would require rich and diverse datasets from a wide range of translations and translation experts. Furthermore, a human-in-the loop [123] approach would incorporate human translators that can give feedback to LLMs so that it can learn and improve from the mistakes in mistranslations. Furthermore, ablation studies can be done to analyse the individual impact of sentiment versus semantic evaluation to review the similarity and difference of the translations by the different LLMs.
@@
Future research could explore improving translations by incorporating contextual background, cultural cues, and domain-specific fine-tuning to enhance both sentiment alignment and semantic accuracy. LLMs offer more advantages over Google Translate and Human Expert since the user can promote LLMs to provide a specific vocabulary style and tone in terms of sentiments. Furthermore, the translated text by human experts can also be transformed by LLMs to get different sets of vocabulary, for example, if the translation of the Bhagavad Gita is from the 1920s, LLMs can be used to give a modern rendering in terms of vocabulary, which would be easier to understand by modern readers. Our framework is general and has been used in previous works where the Bhagavad Gita human expert translation has been compared [65], along with the effectiveness of Google Translate [74]. This study motivates future work to extend the framework for evaluating other languages, particularly high-resource languages, for effective evaluation as more resources and data are available, such as the case of European languages (such as Spanish, French, and Italian). This would better evaluate the cross-cultural strengths and limitations of LLMs in language translation. The framework can also be extended to study how certain languages influence and are coupled with each other, and how certain metaphors and linguistic expressions overlap. Examples include Urdu vs Hindi and Hindi vs Hinglish, which is an evolving language of the internet and prominent in social media platforms such as X (Twitter). In future work, the framework can be extended to review cultural nuances and idiomatic expressions, requiring expert-labelled datasets (such as from Acharya et al. [124]) for the languages under investigation. Furthermore, the framework can be extended to novel LLMs, including multimodal LLMs, so that real-time speech recognition for machine translation can be done. In summary, our study and framework have paved the way for the automated evaluation of language translation models.
Finally, we stress that LLMs must translate sacred texts such as the Bhagavad Gita with caution. This is because the nature of the translations can vary and have different meanings, such as for sensitive topics, such as whether meat eating is prohibited in Hinduism. The subject of this discussion is through the translation of the term “tamasik” diet (Chapter 17, verses 8-10), which is often mistranslated as a meat-eating diet, while “tamas” translates to darkness. Therefore, some commentaries (such as interpretations by the International Society for Krishna Consciousness [125]) take it that the Bhagavad Gita has prohibited meat. Hence, the ethics around the impact of translations for various religious and sacred texts need to be taken into account, and the LLM translation needs to be verified by language experts.
VI Conclusion